Marivic is 20 years older than Fe. Seven years from now, Marivic will be thrice as old as Fe. Find their present ages.
Answer by ptaylor(2198)

Let x=Fe's age; x+7=Fe's age 7 years form now
Then x+20=Marivic's age; x+20+7=x+27=Marivic's age 7 years from now
Now we are told the following:
x+27=3(x+7) get rid of parens (distributive)
x+27=3x+21 subtract x and also 21 from each side
x-x+27-21=3x-x+21-21 collect like terms
6=2x divide each side by 2
x=3----------------------------Fe's age
x+20=3+20=23----------------------Marivic's age
CK
23+7=3(3+7)
30=30
Hope this helps---ptaylor